# Shree Pannalal Patel Marg, Ahmedabad - 380009, Gujarat, India

### v1

@@ 3.2 Permutation invariance and variable cardinality
Restricting events to an ordered and fixed sized representation may result in truncation and combinatorial ambiguities Onyisi:2022hdh . While we may be interested in a specific number of coloured objects at the parton level, due to their highly radiative nature, it is often advantageous to not have a hard restriction on the number of reconstructed jets with additional jets arising from additional radiations. In other words, truncating an event signature to a fixed sized representation introduces ambiguities in how one should select these fixed number of objects with the kinematics inextricably linked to those objects which are not utilised in the representation. Combinatorial ambiguities arise when a signature has more than one way of associating the reconstructed objects to an earlier parton. For instance, in the fully leptonic decay of a pair of top quarks, other then the inability to reconstruct the momentum of the two neutrinos, we also have to decide the allocation of two bottom jets (as we cannot determine the charge) to the two reconstructed lepton. The situation becomes more complicated if we have more than two identified bottom jets in a candidate event.
@@ 3.3 Infrared and collinear safety: resilience to deformations
The presence of massless particles, demand an additional consideration for perturbative calculations in QFT, namely, infrared and collinear (IRC) safety. This is one of the cornerstones for fixed order calculability of any observable for theories containing massless gauge bosons, where there is an intricate cancellation of infrared divergences arising from virtual loops with the interference contribution of soft and collinear emmisions. This cancellation is guaranteed to happen for IRC safe observables, i.e., when an observable defined for $n+1$ particles is equal (or as a limit) to the same for $n$ particles when any of the $n+1$ particles become soft or any two of them become collinear. Therefore, IRC safety puts additional relations between the functions defined on sets with cardinality difference of one.
@@ 3.4 Priors from image processing
Convolutional Neural Networks (CNNs) FUKUSHIMA1982455 ; 6795724 remain the quintessential example of deep learning algorithms with very good generalisation abilities after curtailing the expressivity based on domain knowledge. On top of being approximately resilient to local deformations as explained in the previous subsection, they also assume local connectivity and parameter sharing, which, when combined with succesive composition of layers result in scale separation. These biases are carried over in some sense into graph-based architectures depending on specific details of the architecture which we will touch upon in the next section. Here, we discuss the natural ways in which these three assumptions already account for the structure of QCD radiation patterns arising at the LHC.
@@
Local connectivity means that for a given input pixel or a node, the information processed by the network concentrates in its local neighbourhood. Compared to the image size, the smaller filters in a CNN act on local regions determined by their relative sizes. The collinear splitting structure in parton showers which result in the formation of jets is an inherently local phenomenon in the rapidity-azimuth plane. Therefore, the local nature of feature extraction looks into the local substructure in jets or jets in events. Parameter sharing refers to reusing the same set of weights for different local regions in the input. In QCD, the universality of parton showers results in similar patterns of radiation over the full detector in different positions which motivates reusing the same set of weights to pick up such features. Separation of scales occurs via successive composition of functions acting on local regions in the input representation. Being a more refined version of depth to learn hierarchical abstractions of the data, the locally acting functions create a separation of scales, where the filters in initial layers look at the immediate locality and deeper ones sequentially capture larger scales. This is diagrammatically explained in figure 2. Consequently, hard prongs within jets or jets within events are present at different scales in the rapidity-azimuth plane, and a separation of scales help look into these structures in a hierarchical manner. An example for a three prong top jet is shown in figure 3.
@@ 4.1 Functions on sets
where $\mathbf{h}_{i}$ are the set elements, and $\rho$ and $\phi$ are MLPs. We can see that the function $\phi$ learns a per-object map to a latent space, which undergoes a permution-invariant summation over the set constituents. Going into the two-part seggregation of deep learning algorithms, the MLP $\phi$ acts as a feature extractor while $\rho$ gives the downstream output. The extracted features are combined in a permutation invariant manner under the summation, which reflects the prior of set-based functions. Note that the summation can be replaced by any permutation invariant operation like taking the component-wise maximum or minimum over the latent representation.
@@ 4.2 Describing data with graphs
As we have describe above, graphs equip the input set with relational information between two elements. These elements in graph terminology are called nodes or vertices. The edge set $\mathcal{E}$, is a subset of the cartesian product $\mathcal{S}\times\mathcal{S}$. The presence or absence of an edge, say $(i,j)$ in $\mathcal{E}$ is determined by usefulness of the relation between the two elements $i$ and $j$. Consequently, the way a graph is constructed detemines the information highlighted amongst the nodes. In LHC physics, these may include forming edges between geometrically close components in a detector module, a Feynamn diagram based construction, or closeness in the rapidity-azimuth plane. If $\mathcal{E}=\mathcal{S}\times\mathcal{S}$, the corresponding graph is called a complete graph, the relation between every pair of elements is exposed in the graph. This is generally avoided as it leads to a quadratic scaling with the input set cardinality.
@@
Figure 4: Examples of isomorphic and non-isomorphic graphs between directed graphs with nodes distinguished by their colours. The graphs in the bounded blue box are isomorphic to one another while the three on the righ are not isomorphic to any other graph. The two graphs within the red boxes become isomorphic to the blue ones when one forgoes the difference in the colours and the direction of the edges.
@@ 5 Overview of applications in high energy physics
Graph Neural Networks in their different guises have been explored in a varied range of applications for the physics program at the Large Hadron Collider. There can be several ways to categorise these applications based on the learning algorithm, the prediction task, and the physics application. The categories of learning methods broadly fall under unsupervised, semi-supervised and supervised methods, while the prediction task can be seggregated into node, edge and graph level tasks. In this section, we give a brief overview of the applications of graph-based for some physics applications at the LHC.
@@
As already explained in Section 3.1, physical symmetry has been used in graph-based architectures. Lorentz Group Network (LGN) Bogatskiy:2020tje utilise the isomorphism of proper orthochronous Lorentz group SO$(1,3)^{+}$ to the projective special complex linear group $\text{PSL}(2,\mathbb{C})$ to build tensorial representations decomposed into irreducible representations in terms of the Clebsch-Gordan coefficients. While LorentzNet Gong:2022lye ; Li:2022xfc implemented a simplified scalar and vector representation in the hidden representations to extract the features. Later, PELICAN Bogatskiy:2022czk ; Bogatskiy:2023nnw combined permutation and Lorentz equivariant aggregations. On the public top tagging data set, these architectures matched or outperformed ParticleNet with fewer trainable parameters. Meanwhile, LorentzNet showed a very high sample efficiency, reaching an AUC of 0.9839 with only $5\%$ of the training data. LorentzNet was augmented with capsule network Sahu:2024sts , showing improved performance for quark-gluon performance.
@@
Constructing infrared and collinear-safe feature extractors on point clouds has also received considerable effort. Energy Flow Networks Komiske:2018cqr (EFNs) adapted the Deep SetsNIPS2017_f22e4747 framework to account for IRC safety by utilising per-particle maps of the angular coordinates and performing a summed readout over the particles after linearly weighting with energy. Extensions of this approach include a permutation equivariant EFN Dolan:2020qkr , building hierarchical EFNs by utilising a basis of Legendre polynomials Shen:2023ofd , Lipschitz-EFNs Bright-Thonney:2023gdl where the extracted features are made to follow the Lipschitz continuity condition motivated from the geometry of particle collisions Komiske:2020qhg , and utilising higher moments Gambhir:2024dtf . The per-particle nature of EFN feature extraction has been generalised to graphs in the framework of Energy-weighted Message Passing Networks Konar:2021zdg for supervised classification. Further extensions include hypergraph based feature extraction Konar:2023ptv , and equivariance in the rapidity-azimuth plane Chatterjee:2024pbp ; Bhardwaj:2024djv . An architecture based on dynamic edge convolutions was also studied for tagging jets originating from dark showers Bernreuther:2020vhm .
@@
Another theoretically motivated way to represent a jet as a graph is through the Lund tree declustering Andersson1989 ; Dreyer:2018nbf , which maps emissions according to the jet clustering history. LundNet Dreyer:2020brq employed Lund graphs as inputs for various binary jet classification tasks. Due to the theoretically transparent nature of the input construction, it allows more detailed semi-analytic studies Dreyer:2021hhr of its discriminative power than is possible with the other graph-based approaches.
@@
Graphs can have varying numbers of nodes and edges for each event, unlike fixed-size inputs in typical CNNs. Therefore, GNNs are adaptable to work on different event topologies and can generalize well across various types of particle collisions. In an early study, Abdughani:2018wrw collision events were represented as event graphs, and the message-passing neural network (MPNN) was used to search for the stop pair production at the LHC. Employing GNN for particle-flow (PF) reconstruction in a high-pileup environment has been developed for a general-purpose multilayered particle detector Pata:2021oez . In ref. Anisha:2022ctm , GNN with a fully connected graph is used to combat comprehensive backgrounds arising from other SM processes, which improves the sensitivity of di-Higgs analysis in the vector-boson fusion production channel. Similar architectures are used to improve the four top measurements for SM Anisha:2023xmh and constrain the BSM parameter in the context of 2HDM. The fully connected graph and architecture used in the four top analysis are illustrated in figure 6. Self-attention to extract features from jet substructure and cross attention mechanism to combine these features for event level classification have also been studied Hammad:2023sbd for BSM searches at LHC showing improved performance compared to simple concatenation of features. The triple Higgs boson production bounds are also improved by employing GNN Stylianou:2023xit . Topological graph algorithms were also developed to reconstruct intermediate particles and address combinatorial challenges in full hadronic decays $t\bar{t}$ Ehrke:2023cpn . Recently, hypergraph representations Birch-Sykes:2024gij have been employed to demonstrate the reconstruction of parent particles for full hadronic decays of $t\bar{t}$. Events represented by weighted nodes with edges depicting the distance between events in kinematic space, were utlised Mullin:2019mmh to separate the stop quark pairs from the top pair in the kinematic space. In refs. Pfeffer:2024tjl ; Builtjes:2022usj , $t\bar{t}$ process is scrutinized with Transformer and graph-based event classification. Using the point cloud representation for Higgs boson decays in tau leptons significantly improves compared to classical analysis techniques Onyisi:2022hdh .
@@
A common way to achieve this unsupervised learning is via autoencoders. An autoencoder is an ANN used in unsupervised learning that aims to encode input data into a reduced-dimensional representation (called the latent representation) and then decode it back to its original form. The model learns by comparing the reconstructed output with the original input and tuning parameters to minimize the reconstruction error. GNNs rely on understanding the graph’s topology and node connections to extract meaningful representations. Capturing this information accurately during the encoding and decoding phases to reconstruct the original graph structure is non-trivial. Particle Graph Autoencoder Kasieczka:2021xcg ; Tsan:2021brw used an edge-convolution encoder block to map the node features to a two-dimensional latent node representation, which fed to the symmetrical decoder block to reconstruct the node features. For anomaly detection of events or jets, we are interested in the global features of the graph. Although a graph readout extracts this information in classification, the operation itself is destructive, and one cannot recover the graph structure in an autoencoder-based approach where the decoder needs to have the graph structure intact for its reconstruction. To address these challenges in graph autoencoders, ref, Atkinson:2021nlt devised edge reconstruction networks, which were used to reconstruct weighted edge features, thereby learning the graph structure without undergoing a destructive graph readout operation. The Deep Sets autoencoder Ostdiek:2021bem , on the other hand, relied solely on the encoded global latent space without using any decoder. Graph autoencoders with physics-motivated inductive biases like Lorentz group equivariance Hao:2022zns and IRC-safety Atkinson:2022uzb have also been explored for anomaly detection.
@@ 6 Summary
In this review, we have laid down the rationale for designing automatic feature extractors from high-dimensional data in the context of high-energy physics. The main theme of such strategies is the balance between expressivity and generalisation power. While expressivity deals with the theoretical power of non-parametric models to capture a large set of functions accurately, generalisation power trades some of this expressivity for the practical gain of finding a good function for specific purposes based on priori knowledge of the specific use case.
@@ Acknowledgements
A.B. is supported by the U.S. Department of Energy under grant number DE-SC 0016013. V. S. N. is supported by the STFC under grant ST/P001246/1.

### v2

@@ 3.2 Permutation invariance and variable cardinality
Restricting events to an ordered and fixed sized representation may result in truncation and combinatorial ambiguities Onyisi:2022hdh . While we may be interested in a specific number of coloured objects at the parton level, due to their highly radiative nature, it is often advantageous to not have a hard restriction on the number of reconstructed jets with additional jets arising from additional radiations. In other words, truncating an event signature to a fixed sized representation introduces ambiguities in how one should select these fixed number of objects with the kinematics inextricably linked to those objects which are not utilised in the representation. Combinatorial ambiguities arise when a signature has more than one way of associating the reconstructed objects to an earlier parton. For instance, in the fully leptonic decay of a pair of top quarks, other than the inability to reconstruct the momentum of the two neutrinos, we also have to decide the allocation of two bottom jets (as we cannot determine the charge) to the two reconstructed lepton. The situation becomes more complicated if we have more than two identified bottom jets in a candidate event.
@@ 3.3 Infrared and collinear safety: resilience to deformations
The presence of massless particles, demand an additional consideration for perturbative calculations in QFT, namely, infrared and collinear (IRC) safety. This is one of the cornerstones for fixed order calculability of any observable for theories containing massless gauge bosons, where there is an intricate cancellation of infrared divergences arising from virtual loops with the interference contribution of soft and collinear emissions. This cancellation is guaranteed to happen for IRC safe observables, i.e., when an observable defined for $n+1$ particles is equal (or as a limit) to the same for $n$ particles when any of the $n+1$ particles become soft or any two of them become collinear. Therefore, IRC safety puts additional relations between the functions defined on sets with cardinality difference of one.
@@ 3.4 Priors from image processing
Convolutional Neural Networks (CNNs) FUKUSHIMA1982455 ; 6795724 remain the quintessential example of deep learning algorithms with very good generalisation abilities after curtailing the expressivity based on domain knowledge. On top of being approximately resilient to local deformations as explained in the previous subsection, they also assume local connectivity and parameter sharing, which, when combined with successive composition of layers result in scale separation. These biases are carried over in some sense into graph-based architectures depending on specific details of the architecture which we will touch upon in the next section. Here, we discuss the natural ways in which these three assumptions already account for the structure of QCD radiation patterns arising at the LHC.
@@
Local connectivity means that for a given input pixel or a node, the information processed by the network concentrates in its local neighbourhood. Compared to the image size, the smaller filters in a CNN act on local regions determined by their relative sizes. The collinear splitting structure in parton showers which result in the formation of jets is an inherently local phenomenon in the rapidity-azimuth plane. Therefore, the local nature of feature extraction looks into the local substructure in jets or jets in events. Parameter sharing refers to reusing the same set of weights for different local regions in the input. In QCD, the universality of parton showers results in similar patterns of radiation over the full detector [^10] Even though pile-up break this generality, this remains true after pileup removal after which phenomenological analyses are conducted on events following such patterns. in different positions which motivates reusing the same set of weights to pick up such features. Separation of scales occurs via successive composition of functions acting on local regions in the input representation. Being a more refined version of depth to learn hierarchical abstractions of the data, the locally acting functions create a separation of scales, where the filters in initial layers look at the immediate locality and deeper ones sequentially capture larger scales. This is diagrammatically explained in figure 2. Consequently, hard prongs within jets or jets within events are present at different scales in the rapidity-azimuth plane, and a separation of scales help look into these structures in a hierarchical manner. An example for a three prong top jet is shown in figure 3.
@@ 4.1 Functions on sets
where $\mathbf{h}_{i}$ are the set elements, $M$ the cardinality of the set, and $\rho$ and $\phi$ are MLPs. We can see that the function $\phi$ learns a per-object map to a latent space, which undergoes a permutation-invariant summation over the set constituents. Going into the two-part segregation of deep learning algorithms, the MLP $\phi$ acts as a feature extractor while $\rho$ gives the downstream output. The extracted features are combined in a permutation invariant manner under the summation, which reflects the prior of set-based functions. Note that the summation can be replaced by any permutation invariant operation like taking the component-wise maximum or minimum over the latent representation.
@@ 4.2 Describing data with graphs
As we have describe above, graphs equip the input set with relational information between two elements. These elements in graph terminology are called nodes or vertices. The edge set $\mathcal{E}$, is a subset of the Cartesian product $\mathcal{S}\times\mathcal{S}$. The presence or absence of an edge, say $(i,j)$ in $\mathcal{E}$ is determined by usefulness of the relation between the two elements $i$ and $j$. Consequently, the way a graph is constructed determines the information highlighted amongst the nodes. In LHC physics, these may include forming edges between geometrically close components in a detector module, a Feynman diagram based construction, or closeness in the rapidity-azimuth plane. If $\mathcal{E}=\mathcal{S}\times\mathcal{S}$, the corresponding graph is called a complete graph, the relation between every pair of elements is exposed in the graph. This is generally avoided as it leads to a quadratic scaling with the input set cardinality.
@@
Figure 4: Examples of isomorphic and non-isomorphic graphs between directed graphs with nodes distinguished by their colours. The graphs in the bounded blue box are isomorphic to one another while the three on the right are not isomorphic to any other graph. The two graphs within the red boxes become isomorphic to the blue ones when one forgoes the difference in the colours and the direction of the edges.
@@ 5 Overview of applications in high energy physics
Graph Neural Networks in their different guises have been explored in a varied range of applications for the physics program at the Large Hadron Collider. There can be several ways to categorise these applications based on the learning algorithm, the prediction task, and the physics application. The categories of learning methods broadly fall under unsupervised, semi-supervised and supervised methods, while the prediction task can be segregated into node, edge and graph level tasks. In this section, we give a brief overview of the applications of graph-based for some physics applications at the LHC.
@@
As already explained in Section 3.1, physical symmetry has been used in graph-based architectures. Lorentz Group Network (LGN) Bogatskiy:2020tje utilise the isomorphism of proper orthochronous Lorentz group SO$(1,3)^{+}$ to the projective special complex linear group $\text{PSL}(2,\mathbb{C})$ to build tensorial representations decomposed into irreducible representations in terms of the Clebsch-Gordan coefficients. While LorentzNet Gong:2022lye ; Li:2022xfc implemented a simplified scalar and vector representation in the hidden representations to extract the features. Later, PELICAN Bogatskiy:2022czk ; Bogatskiy:2023nnw combined permutation and Lorentz equivariant aggregations. On the public top tagging data set, these architectures matched or outperformed ParticleNet with fewer trainable parameters. Meanwhile, LorentzNet showed a very high sample efficiency, reaching an AUC of 0.9839 with only $5\%$ of the training data. LorentzNet was augmented with capsule network Sahu:2024sts , showing improved performance for quark-gluon performance. The gain in performance comes because the feature extraction specifically utilises the physical symmetry of the dataset.
@@
When utilising the jet’s constituents when extracting features with deep-learning models, it is not necessary that automatic feature extraction on a jet is IRC safe, and the IRC safety of these features depend not only on the input features but also on how they are structured. Constructing infrared and collinear-safe feature extractors on point clouds has also received considerable effort. Energy Flow Networks Komiske:2018cqr (EFNs) adapted the Deep SetsNIPS2017_f22e4747 framework to account for IRC safety by utilising per-particle maps of the angular coordinates and performing a summed readout over the particles after linearly weighting with energy. Extensions of this approach include a permutation equivariant EFN Dolan:2020qkr , building hierarchical EFNs by utilising a basis of Legendre polynomials Shen:2023ofd , Lipschitz-EFNs Bright-Thonney:2023gdl where the extracted features are made to follow the Lipschitz continuity condition motivated from the geometry of particle collisions Komiske:2020qhg , and utilising higher moments Gambhir:2024dtf . The per-particle nature of EFN feature extraction has been generalised to graphs in the framework of Energy-weighted Message Passing Networks Konar:2021zdg for supervised classification. Further extensions include hypergraph based feature extraction Konar:2023ptv , and equivariance in the rapidity-azimuth plane Chatterjee:2024pbp ; Bhardwaj:2024djv . An architecture based on dynamic edge convolutions was also studied for tagging jets originating from dark showers Bernreuther:2020vhm .
@@
Another theoretically motivated way to represent a jet as a graph is through the Lund tree de-clustering Andersson1989 ; Dreyer:2018nbf , which maps emissions according to the jet clustering history. LundNet Dreyer:2020brq employed Lund graphs as inputs for various binary jet classification tasks. Due to the theoretically transparent nature of the input construction, it allows more detailed semi-analytic studies Dreyer:2021hhr of its discriminative power than is possible with the other graph-based approaches.
@@
Graphs can have varying numbers of nodes and edges for each event, unlike fixed-size inputs in typical CNNs. Therefore, GNNs are adaptable to work on different event topologies and can generalize well across various types of particle collisions. In an early study, Abdughani:2018wrw collision events were represented as event graphs, and the message-passing neural network (MPNN) was used to search for the stop pair production at the LHC. Employing GNN for particle-flow (PF) reconstruction in a high-pileup environment has been developed for a general-purpose multilayered particle detector Pata:2021oez . In ref. Anisha:2022ctm , GNN with a fully connected graph is used to combat comprehensive backgrounds arising from other SM processes, which improves the sensitivity of di-Higgs analysis in the vector-boson fusion production channel. Similar architectures are used to improve the four top measurements for SM Anisha:2023xmh and constrain the BSM parameter in the context of 2HDM. The fully connected graph and architecture used in the four top analysis are illustrated in figure 6. Self-attention to extract features from jet substructure and cross attention mechanism to combine these features for event level classification have also been studied Hammad:2023sbd for BSM searches at LHC showing improved performance compared to simple concatenation of features. Accurate measurement of the trilinear and quartic Higgs self-couplings is essential to understanding the shape of the Higgs potential and the electroweak phase transition. While traditional searches at the LHC using Higgs pair production provide limited constraints, recent studies using Graph Neural Networks (GNNs) have shown promise in improving these bounds Stylianou:2023xit . Topological graph algorithms were also developed to reconstruct intermediate particles and address combinatorial challenges in full hadronic decays $t\bar{t}$ Ehrke:2023cpn . Recently, hypergraph representations Birch-Sykes:2024gij have been employed to demonstrate the reconstruction of parent particles for full hadronic decays of $t\bar{t}$. Events represented by weighted nodes with edges depicting the distance between events in kinematic space, were utilised Mullin:2019mmh to separate the stop quark pairs from the top pair in the kinematic space. In refs. Pfeffer:2024tjl ; Builtjes:2022usj , $t\bar{t}$ process is scrutinized with Transformer and graph-based event classification. Using the point cloud representation for Higgs boson decays in tau leptons significantly improves compared to classical analysis techniques Onyisi:2022hdh .
@@
A common way to achieve this unsupervised learning is via autoencoders. An autoencoder is an ANN used in unsupervised learning that aims to encode input data into a reduced-dimensional representation (called the latent representation) and then decode it back to its original form. The model learns by comparing the reconstructed output with the original input and tuning parameters to minimize the reconstruction error. Potential of autoencoder architectures have been demonstrated in simulating realistic and diverse aspects of high energy physics events. For instance, convolutional autoencoders were studied for detecting anomalous jet images in refs.  Heimel:2018mkt ; Farina:2018fyg .
GNNs rely on understanding the graph’s topology and node connections to extract meaningful representations. Capturing this information accurately during the encoding and decoding phases to reconstruct the original graph structure is non-trivial. Particle Graph Autoencoder Kasieczka:2021xcg ; Tsan:2021brw used an edge-convolution encoder block to map the node features to a two-dimensional latent node representation, which fed to the symmetrical decoder block to reconstruct the node features. For anomaly detection of events or jets, we are interested in the global features of the graph. Although a graph readout extracts this information in classification, the operation itself is destructive, and one cannot recover the graph structure in an autoencoder based approach where the decoder needs to have the graph structure intact for its reconstruction. To address these challenges in graph autoencoders, ref. Atkinson:2021nlt devised edge reconstruction networks, which were used to reconstruct weighted edge features, thereby learning the graph structure without undergoing a destructive graph readout operation. The Deep Sets autoencoder Ostdiek:2021bem , on the other hand, relied solely on the encoded global latent space without using any decoder. Graph autoencoders with physics-motivated inductive biases like Lorentz group equivariance Hao:2022zns and IRC-safety Atkinson:2022uzb have also been explored for anomaly detection.
@@ 6 Summary
In this review, we have laid down the rationale for designing automatic feature extractors from high-dimensional data in the context of high-energy physics. The main theme of such strategies is the balance between expressivity and generalisation power. While expressivity deals with the theoretical power of non-parametric models to capture a large set of functions accurately, generalisation power trades some of this expressivity for the practical gain of finding a good function for specific purposes based on prior knowledge of the use case.
@@ Acknowledgements
A.B. is supported by the U.S. Department of Energy under grant number DE-SC 0016013. V. S. N. is supported by the STFC under grant ST/P001246/1. The computational works are performed using the Param Vikram-1000 High-Performance Computing Cluster and the TDP project resources of the Physical Research Laboratory (PRL). Authors would like to thank for the warm hospitality and support received at the school of Deep Machine Learning for particle and astroparticle physics (ML4HEP) at ICTS, Bengaluru (2023) and IOP, Bhubaneswar (2024).
Data Availability Statement: No data associated in the manuscript.